# Data-driven inventory management for new products: A warm-start and adjusted Dyna-$Q$ approach

### v2

@@ Abstract
In this paper, we propose a novel reinforcement learning algorithm for inventory management of newly launched products with no or limited historical demand information. The algorithm follows the classic Dyna-$Q$ structure, balancing the model-based and model-free approaches, while accelerating the training process of Dyna-$Q$ and mitigating the model discrepancy generated by the model-based feedback. Warm-start information from the demand data of existing similar products can be incorporated into the algorithm to further stabilize the early-stage training and reduce the variance of the estimated optimal policy. Our approach is validated through a case study of bakery inventory management with real data. The adjusted Dyna-$Q$ shows up to a 23.7% reduction in average daily cost compared with $Q$-learning, and up to a 77.5% reduction in training time within the same horizon compared with classic Dyna-$Q$. By incorporating the warm-start information, it can be found that the adjusted Dyna-$Q$ has the lowest total cost, lowest variance in total cost, and relatively low shortage percentages among all the algorithms under a 30-day testing.
@@ I INTRODUCTION
Inventory management is crucial for supply chain operations, overseeing and controlling the order, storage, and usage of goods in a business [1]. In inventory management, the cold-start setting refers to predicting demand and formulating appropriate inventory strategies when new products are introduced or new market demands arise due to the lack of historical data [2]. For cold-start inventory management, data-driven flexible inventory management strategies are adopted, such as Just-in-Time (JIT) inventory management [3], inventory optimization models [4], and predictive analytics and intelligent methods [5].
@@
Reinforcement learning (RL) has been increasingly applied in inventory management to optimize inventory control policies and decision-making processes [5]. Most research on RL’s application in inventory management mainly addresses the problem with existing similar products, whose historical sales information is accessible [6]. Under this scenario, the RL method’s agent has adequate data to obtain a satisfactory ordering strategy [7]. Most studies apply model-free RL in inventory control under sufficient data due to model-free RL’s satisfactory performance under stochastic conditions [8]. On the other hand, model-based RL has the potential to improve upon model-free methods by offering more stable and interpretable policies and potentially faster convergence. However, for newly launched products without actual demand data, the demand prediction model is trained only based on the historical demands of existing products [7]. The discrepancy between the trained model and the actual environment may prevent the model-based algorithm from converging toward the near-optimal policy [9]. Consequently, incorporating demand information from the actual environment obtained by model-free RL methods could benefit the model-based algorithm by obtaining a better result. Therefore, both model-free RL and model-based RL have demonstrated significant potential in optimizing inventory management policies and decision-making processes. However, challenges persist in how to merge their merits together in inventory management for new products with no or limited historical demand data.
@@
The Dyna-$Q$ algorithm combines model-based RL with model-free RL. In the cold-start inventory management problem, the Dyna-$Q$ algorithm learns from both a continuously updated model and the real environment to update the value function or policy function [10]. This approach helps address inventory forecasting and management challenges in situations where historical data is not available. By utilizing simulation and planning, Dyna-$Q$ compensates for the absence of actual data, allowing it to learn and adapt to new environments incrementally. Simulating an environment model that accurately reflects the dynamics of the real environment is challenging, especially in the early stages of training. If the model deviates from the actual environment, policy updates based on this simulated data may lead to suboptimal results. Even in cases where a relatively accurate model can be simulated, maintaining this model demands significant computational resources and data, particularly in complex or dynamically changing environments [11, 12].
@@
To address the existing limitations, this study introduces a warm-start and adjusted Dyna-$Q$ approach. Firstly, the adjusted Dyna-$Q$ algorithm is proposed, which incorporates the search-then-convergence (STC) process to design the iteration-decaying exploration and planning steps, dynamically reweighting the model-based and model-free components. Secondly, a demand forecasting model is constructed based on the demand for existing similar products to provide warm-start prior knowledge of the new product’s demand distribution for the adjusted Dyna-$Q$ algorithm. Finally, the proposed approach is validated using real data from a bakery to manage inventory for a newly launched product. The results demonstrate the superior performance of the proposed approach in terms of cost and variance reduction compared to classic Dyna-$Q$ and $Q$-learning algorithms.
@@ II Framework of the Proposed Approach
This section introduces the proposed algorithm, which is based on the classic Dyna-$Q$ method and aims to minimize long-term discounted operation costs by determining the optimal ordering quantity when the demand distribution of a new product is unknown. The proposed algorithm offers two main innovations. Firstly, the selection probability of arbitrary actions in the epsilon-greedy strategy and the number of planning steps are determined using the search-then-convergence (STC) process [13], which decays with the iteration over the training horizons. This approach accelerates the training process and places greater emphasis on exploiting the model-free components as more experience is gathered from the environment. Secondly, since there is limited historical demand data for the new product during the early training phases, the proposed algorithm incorporates the model-based aspect by utilizing simulated data from existing similar products as warm-start information. These two innovations will be further elaborated in the subsequent subsections.
Denote state (the inventory levels) space $\mathbb{S}$ and for any $s\in\mathbb{S}$, an action (the ordering quantities) space $\mathbb{A}(s)$. Assume that the state and action spaces are all finite. Algorithm 1 follows the structure of Dyna-$Q$, which combines both model-free and model-based methods. In each iteration $t$ within the training horizons, the algorithm directly interacts with the environment and updates the $Q$-values following model-free $Q$-learning. For a given state-action pair $s\in\mathbb{S},\,a\in\mathbb{A}(s)$,
@@
where $c$ denotes the immediate cost, $s^{\prime}\in\mathbb{S}$ the next state transited, $10<\alpha<$ 1 the learning rate, and $10<\gamma<$ 1 the discount factor. The new experience $c$ and $s^{\prime}$ are used in updating the model. Here we use $M(s,\,a)$ to denote the model associated with state-action pair $(s,\,a)$. Explicitly, the model $M(s,\,a)$ consists of two neural networks estimating the state transition probabilities $P(s^{\prime}|s,\,a)$ and the cost function $C(s,\,a,\,s^{\prime})$, respectively. The estimation is updated by adjusting the parameters to minimize the least squares error between the predicted state transition and cost and the new experience [14].
In terms of the model-based part, multiple planning steps are conducted based on the model simulation, where the $Q$-table is further updated, using the same rule as formulation (1).
@@
II-A Adjusted Exploration Strategy and Planning Steps
@@
In each iteration, an epsilon-greedy strategy has been implemented to balance the exploitation and exploration processes. In Algorithm 1, we adjust $\epsilon_{t}$ based on the search-then-convergence (STC) process. [13] In the early iterations of training, the agent tends to explore more rather than exploit the current best action. As iteration evolves, the epsilon value will decrease to a specific value, and the agent will emphasize more on exploitation. The epsilon value is updated based on the STC procedure as below:
$\varepsilon_{t}=\max\left\{\frac{\varepsilon_{0}}{1+y},\,\varepsilon_{\min}\right\},$ (2)
where
$y=\frac{t^{2}}{\partial_{\varepsilon}+t}.$ (3)
Here $\varepsilon_{0}>0$ represents the initial value of the epsilon, $\varepsilon_{\min}>0$ represents the lower bound of the epsilon, and $\partial_{\varepsilon}$ is a constant smoothing the decaying process, which will be chosen appropriately in practice.
@@
The planning component (model-based part) can help the agent examine and update the $Q$-table by exploiting past experiences from the environment. It is known for reducing the sample complexity, leading to faster convergence, and variance reduction [15][16]. However, a trade-off exists as it is less computationally efficient than model-free $Q$-learning within the same training horizons. Algorithm 1 enables acceleration by setting the total planning steps decaying with the iteration, which means that the model-free part is more emphasized as the iteration evolves. The intention is that in the late training horizons, the agent has obtained enough exploration for the environment and can rely on model-free parts, which is computationally efficient. The planning steps $N_{t}$ are updated based on the STC procedure as well:
$N_{t}=\max\left\{\frac{N_{0}}{1+y},\,N_{\min}\right\},$ (4)
where
$y=\frac{t^{2}}{\partial_{N}+t}.$ (5)
Here, $N_{0}>0$ represents the initial value of the planning steps, $N_{\min}>0$ represents the lower bound of the planning steps, and $\partial_{N}$ is a constant smoothing the decaying process, which will be properly chosen in practice.
@@
The anticipated benefit of the adjusted Dyna-$Q$ algorithm is that it accelerates the training process while achieving lower costs in both training and testing phases.
@@
  Update epsilon $\epsilon_{t}$ and planning steps $N_{t}$ based on STC process by equations (2) (3) and (4)(5), respectively.
@@
II-B Warm-start Via Forecasting and Simulation
The proposed Algorithm 1 operates as a cold-start algorithm without prior demand information for new products. This absence of demand information at the beginning can cause discrepancies between the estimated model and the actual environment in the early training phases, potentially resulting in suboptimal policy estimates. To address this issue, a warm-start version of the algorithm is suggested, integrating a simulation-based demand forecasting model to provide advance knowledge of new product demand for the adjusted Dyna-$Q$ algorithm. The demand forecasting model is structured as a three-layer Bayesian neural network (BNN), trained on historical data of similar existing products. BNN can leverage prior knowledge, such as product sales seasonality and market trends, to offer improved predictions of new product demand [17]. This demand forecasting model, denoted as $f_{\theta}(u,v)$, utilizes numerical values related to historical demands and moving averages within a specific window size denoted by $u$, along with multiple categorical features related to calendar data for the day before the demand, marked as $v$. These features include day of the week, weekday/weekend indicator, week number, and sine and cosine values representing the days elapsed in the month and year. Here, $\theta$ represents the model parameters. By leveraging the demand forecasting model, an offline simulated demand dataset $D=[\hat{d}_{1},...,\hat{d}_{h}]$ can be generated by inputting the corresponding day features, with $h$ representing the prediction horizon.
The warm-start information is generated as follows. Before implementing Algorithm 1, we deploy $Q$-learning on the offline data set $D$ and obtain an updated $Q$-table, denoted as $Q_{D}$, as the initialized $Q$-table in Algorithm 1. By employing $Q$-learning, the model is updated using the offline dataset $D$ and referred to as $M_{D}$. This updated model $M_{D}$ serves as the initialization for Algorithm 1. Fig.1 provides a concise overview of the structure of the warm-start adjusted Dyna-$Q$ approach. Incorporating this warm-start information into Algorithm 1 allows the simulated demand from similar existing products to stabilize the demand distribution shift of the new product in the early training phases, thereby reducing training variance. As more new experiences are gathered, the discrepancy between the demand distributions of existing and new products can be mitigated. The adjusted approach introduced in Section II.$A$ can further aid in minimizing this discrepancy.
Figure 1: Pipeline of warm-start and adjusted Dyna-$Q$
@@ III Case Study
In this case study, we are applying the proposed algorithms to address a real-life single-echelon inventory problem faced by a bakery. The bakery store conducts periodic inventory reviews and places orders from a single, reliable supplier with an unlimited supply capacity. The inventory review and reorder period are set to one day, with the assumption that the bakery receives the ordered products immediately. Given the perishable nature of the products, it is assumed that the quality decay process begins upon immediate receipt by the store and lasts for three days. Any unsold products held in inventory for three days will be discarded. The bakery has introduced a new product called Boule 400g (Boule, from French, meaning ”ball”, is a traditional shape of French bread resembling a squashed ball), with its daily demand assumed to follow a random and identically distributed pattern. A dataset containing daily transaction details of customers from a bakery in France, collected from Kaggle [18], provides the daily transaction information from 2021-01-01 to 2022-09-3 for the existing similar product: Boule 200g. The historical demand data for Boule 200g from the bakery can be leveraged to generate warm-start information for the new product, Boule 400g.
The sequence of the bakery’s behaviors is elaborated as follows. At the beginning of each day, the store receives the orders placed yesterday, and the inventory level is updated. Then, the products are sold on a FIFO (First-in and first-out) basis and unfulfilled demand is considered a loss of sales. Next, the quality of products is monitored, and the products lasting in inventory for three days will be discarded. Finally, the inventory level is updated again, and the store places replenishment orders. We formulate the problem and environment in the following subsection.

### v3

# Data-driven inventory management for new products: An adjusted Dyna-$Q$ approach with transfer learning

@@ Abstract
In this paper, we propose a novel reinforcement learning algorithm for inventory management of newly launched products with no historical demand information. The algorithm follows the classic Dyna-$Q$ structure, balancing the model-free and model-based approaches, while accelerating the training process of Dyna-$Q$ and mitigating the model discrepancy generated by the model-based feedback. Based on the idea of transfer learning, warm-start information from the demand data of existing similar products can be incorporated into the algorithm to further stabilize the early-stage training and reduce the variance of the estimated optimal policy. Our approach is validated through a case study of bakery inventory management with real data. The adjusted Dyna-$Q$ shows up to a 23.7% reduction in average daily cost compared with $Q$-learning, and up to a 77.5% reduction in training time within the same horizon compared with classic Dyna-$Q$. By using transfer learning, it can be found that the adjusted Dyna-$Q$ has the lowest total cost, lowest variance in total cost, and relatively low shortage percentages among all the benchmarking algorithms under a 30-day testing.
@@ I INTRODUCTION
Inventory management is crucial for supply chain operations, overseeing and controlling the order, storage, and usage of goods in business [1]. In inventory management, the cold-start setting refers to predicting demand and formulating appropriate inventory strategies when new products are introduced or new market demands arise due to the lack of historical data [2]. Data-driven flexible inventory management strategies are adopted, such as Just-in-Time inventory management [3], inventory optimization models [4], and predictive analytics and intelligent methods [5].
@@
Reinforcement learning (RL) has been increasingly applied in inventory management to optimize inventory control policies and decision-making processes [5]. Most research on RL’s application in inventory management mainly addresses the problems with existing similar products, whose historical sales information is accessible [6]. The RL method’s agent has adequate data to obtain a satisfactory ordering strategy [7]. Most studies apply model-free RL in inventory control under sufficient data due to model-free RL’s satisfactory performance under stochastic conditions [8]. On the other hand, model-based RL has the potential to improve upon model-free methods by offering more stable and interpretable policies and potentially faster convergence. However, for newly launched products without actual demand data, the demand prediction model is trained only based on the historical demands of existing products [7]. The discrepancy between the trained model and the actual environment may prevent the model-based algorithm from converging toward the near-optimal policy [9]. Consequently, incorporating demand information from the actual environment obtained by model-free RL methods could benefit the model-based algorithm by obtaining a better result. Therefore, both model-free RL and model-based RL have demonstrated significant potential in optimizing inventory management policies and decision-making processes. However, challenges persist in how to merge their merits together in inventory management for new products with no historical demand data.
@@
The Dyna-$Q$ algorithm combines model-based RL with model-free RL, which learns from a continuously updated model and the real environment, to update the value function or policy function [10]. This approach helps address inventory forecasting and management challenges when historical data is unavailable. By utilizing simulation and planning, Dyna-$Q$ compensates for the absence of actual data, allowing it to learn and adapt to new environments incrementally. Simulating an environment model that accurately reflects the dynamics of the real environment is challenging, especially in the early stages of training. If the model deviates from the actual environment, policy updates based on the simulated data may lead to suboptimal results. Even in cases where a relatively accurate model can be simulated, maintaining this model demands significant computational resources and data, particularly in complex or dynamically changing environments [11, 12].
There exist technologies to enhance learning efficiency and performance through the application of knowledge acquired from one task to another related task, which constitutes a form of transfer learning. Transfer $Q$-learning frequently expedites the learning process for a new task by transferring knowledge of models, $Q$-tables, targets, features, and parameters from a source task [13]. An Imitation and Transfer $Q$-learning (ITQ) approach has been proposed for load parameter identification problem [14], which transfers the $Q$-value from the source task as the initial start of the new task. ITQ effectively balances greedy and global search strategies, thereby preventing premature convergence and attaining global optimum, with simulation results demonstrating superior convergence compared to traditional methodologies. Target transfer $Q$-learning has been introduced to transfer the $Q$-function from a source task as the temporal difference target for a new task under safe conditions, ensuring more rapid convergence than standard $Q$-learning [15]. Comprehensive task models were transferred in hybrid RL agents utilizing inter-task mappings and cascade neural networks, amalgamating model-free and model-based learning to achieve substantial performance enhancements over non-transfer approaches [16]. Empirical results confirm its 43.58% improvement in efficiency relative to conventional deep transfer learning methods, representing the inaugural study to optimize knowledge transfer for deep learning in networking applications [17]. To the best of our knowledge, no research has been done that integrates transfer learning with Dyna-$Q$ within the context of inventory control problems. There is also little work on transferring both $Q$-value and the model of a source task in model-based RL.
@@
To address the existing limitations mentioned above, this study firstly introduces an adjusted Dyna-$Q$ algorithm, which incorporates the search-then-convergence (STC) process to design the iteration-decaying exploration and planning steps, dynamically reweighting the model-based and model-free components. Secondly, based on the idea of transfer learning, a demand forecasting model is constructed based on the demand for existing similar products to provide warm-start prior knowledge of the new product’s demand distribution for the adjusted Dyna-$Q$ algorithm. Finally, the proposed approach is validated using real data from a bakery to manage inventory for a newly launched product. The results demonstrate the superior performance of the proposed approach in terms of cost and variance reduction compared to classic Dyna-$Q$ and $Q$-learning algorithms.
@@ II Framework of the Proposed Approach
This section introduces the proposed algorithm, which is developed based on the classic Dyna-$Q$ method and aims to minimize long-term discounted operation costs by determining the optimal ordering quantity when the demand distribution of a new product is unknown. The proposed algorithm offers two main innovations. Firstly, the selection probability of arbitrary actions in the epsilon-greedy strategy and the number of planning steps are determined using the search-then-convergence (STC) process [18], which decays with the iteration over the training horizons. This approach accelerates the training process and places greater emphasis on exploiting the model-free components as more experience is gathered from the environment. Secondly, since there is no historical demand data for the new product during the early training phases, the proposed algorithm incorporates transfer learning by utilizing simulated data from existing similar products as warm-start information. These two innovations will be further elaborated in the subsequent subsections.
Denote state (the inventory levels) space $\mathbb{S}$ and for any $s\in\mathbb{S}$, an action (the ordering quantities) space $\mathbb{A}(s)$. Assume that the state and action spaces are all finite. The proposed Algorithm 1 follows the structure of Dyna-$Q$, which combines both model-free and model-based methods. In each iteration $t$ within the training horizons, the algorithm directly interacts with the environment and updates the $Q$-values following model-free $Q$-learning. For a given state-action pair $s\in\mathbb{S},\,a\in\mathbb{A}(s)$,
@@
where $c$ denotes the immediate cost, $s^{\prime}\in\mathbb{S}$ the next state transited, $10<\alpha<$ 1 the learning rate, and $10<\gamma<$ 1 the discount factor. The new experience $c$ and $s^{\prime}$ are used in updating the model. Here, we use $M(s,\,a)$ to denote the model associated with state-action pair $(s,\,a)$. Explicitly, the model $M(s,\,a)$ consists of two neural networks estimating the state transition probabilities $P(s^{\prime}|s,\,a)$ and the cost function $C(s,\,a,\,s^{\prime})$, respectively. The estimation is updated by adjusting the parameters to minimize the least squares error between the predicted state transition and cost and the new experience [19]. In terms of the model-based part, multiple planning steps are conducted based on the model simulation, where the $Q$-table is further updated, using the same rule as formulation (1).
@@
II-A Adjusted Exploration Strategy and Planning Steps based on Search-then-Convergence (STC) Process
@@
An epsilon-greedy strategy has been implemented in each iteration to balance the exploitation and exploration processes. In Algorithm 1, we adjust $\epsilon_{t}$ based on the search-then-convergence (STC) process [18]. In the early iterations of training, the agent tends to explore more rather than exploit the current best action. As iteration evolves, the epsilon value will decrease to a specific value, and the agent will emphasize more on exploitation. The epsilon value is updated based on the STC procedure as below: $\varepsilon_{t}=\max\left\{\varepsilon_{0}/1+y,\,\varepsilon_{\min}\right\}$, where $y=t^{2}/(\partial_{\varepsilon}+t)$. Here $\varepsilon_{0}>0$ represents the initial value of the epsilon, $\varepsilon_{\min}>0$ represents the lower bound of the epsilon, and $\partial_{\varepsilon}$ is a constant smoothing the decaying process, which will be chosen appropriately in practice.
@@
The planning component (model-based part) can help the agent examine and update the $Q$-table by exploiting past experiences from the environment. It is known for reducing the sample complexity, leading to faster convergence and variance reduction [20][21]. However, a trade-off exists as it is less computationally efficient than model-free $Q$-learning within the same training horizons. Algorithm 1 enables acceleration by setting the total planning steps decaying with the iteration, which means that the model-free part is more emphasized as the iteration evolves. The intention is that in the late training horizons, the agent has obtained enough exploration for the environment and can rely on model-free part, which is computationally efficient. The planning steps $N_{t}$ are updated based on the STC procedure as well: $N_{t}=\max\left\{N_{0}/1+y,\,N_{\min}\right\},$ where $y=t^{2}/(\partial_{N}+t)$. Here, $N_{0}>0$ represents the initial value of the planning steps, $N_{\min}>0$ represents the lower bound of the planning steps, and $\partial_{N}$ is a constant smoothing the decaying process, which will be chosen appropriately in practice.
@@
The anticipated benefit of the adjusted Dyna-$Q$ algorithm is that it accelerates the training process while achieving lower costs in both the training and testing phases.
@@
  Update epsilon $\epsilon_{t}$ and planning steps $N_{t}$ based on STC process.
@@
II-B Warm-start Information from Transfer Learning
Algorithm 1 operates as a cold-start algorithm without prior demand information for new products. This absence of demand information at the beginning can cause discrepancies between the estimated model and the actual environment in the early training phases, potentially resulting in suboptimal policy estimates. To address this issue, another advanced variant of the algorithm with transfer learning is developed, integrating a simulation-based demand forecasting model to provide warm-start knowledge of new product demand for the adjusted Dyna-$Q$ algorithm. The demand forecasting model is structured as a three-layer Bayesian neural network (BNN), trained on historical data of similar existing products. BNN can leverage prior knowledge, such as product sales seasonality and market trends, to offer improved predictions of new product demand [22]. This demand forecasting model, denoted as $f_{\theta}(u,v)$, utilizes numerical values related to historical demands and moving averages within a specific window size denoted by $u$, along with multiple categorical features related to calendar data for the day before the demand, marked as $v$. These features include: day of the week, weekday/weekend indicator, week number, and sine and cosine values representing the days elapsed in the month and year. Here, $\theta$ represents the model parameters. By leveraging the demand forecasting model, an offline simulated demand dataset $D=[\hat{d}_{1},...,\hat{d}_{h}]$ can be generated by inputting the corresponding day features, with $h$ representing the prediction horizon.
The transfer learning process is executed as follows. Before implementing Algorithm 1, we deploy $Q$-learning on the offline data set $D$ and obtain an updated $Q$-table, denoted as $Q_{D}$, which will be transferred as the initial $Q$-table in Algorithm 1. The updated model $M_{D}$ is transferred as the initialization for Algorithm 1. Figure 1 provides a concise overview of the structure of the adjusted Dyna-$Q$ approach with transfer learning. Incorporating transfer learning into Algorithm 1 allows the simulated demand data from similar existing products to stabilize the demand distribution shift of the new product in the early training phases, thereby reducing training variance. As more new experiences are gathered, the discrepancy between the demand distributions of existing and new products can be mitigated. The adjusted approach introduced in Section II.$A$ can further aid in minimizing this discrepancy.
Figure 1: Pipeline of adjusted Dyna-$Q$ with transfer learning
@@ III Case Study
In this case study, we apply the proposed algorithms to address a real-life single-echelon inventory problem faced by a bakery. The bakery store conducts periodic inventory reviews and places orders from a single, reliable supplier with an unlimited supply capacity. The inventory review and reorder period are set to one day, assuming that the bakery receives the ordered products immediately. Given the perishable nature of the products, it is assumed that the quality decay process begins upon immediate receipt by the store and lasts for three days. Any unsold products held in inventory for three days will be discarded. The bakery has introduced a new product called Boule 400g (Boule, from French, meaning ”ball”, is a traditional shape of French bread resembling a squashed ball), with its daily demand assumed to follow a random and identically distributed pattern. A dataset containing daily transaction details of customers from a bakery in France, collected from Kaggle [23], provides the daily transaction information from 2021-01-01 to 2022-09-3 for the existing similar product: Boule 200g. The historical demand data for Boule 200g from the bakery can be leveraged to generate warm-start information for the new product, Boule 400g.
The sequence of the bakery’s behaviors is elaborated as follows. At the beginning of each day, the store receives the orders placed yesterday, and the inventory level is updated. Then, the products are sold on a FIFO (First-in and first-out) rule and unfulfilled demand is considered a loss of sales. Next, the quality of products is monitored, and the products lasting in inventory for three days will be discarded. Finally, the inventory level is updated again, and the store places replenishment orders. We formulate the problem and environment in the following subsection.